OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 5 is 3 more than –6. What is the number? A. –45 B. –30 C. –15 D. –5

OpenStudy (texaschic101):

x/5 = -6 + 3 can you solve this ?

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

ohhh is it A?

OpenStudy (texaschic101):

nope... x/5 = -6 + 3 x/5 = -3 --- multiply both sides by 5 x = (5)(-3) x = -15 you see, if you multiply x/5 by 5, the 5's cancel leaving you with just x...but what you do on one side of the equal sign, you have to do on the other...so then you multiply -3 by 5 giving you : x = -15 any questions ?

OpenStudy (anonymous):

oh, i didnt see the multiply by 5 part

OpenStudy (anonymous):

Thank you :)

OpenStudy (texaschic101):

anytime....just remember, if you have a fraction in the problem, whether it be an equation or an inequality, if you multiply the entire equation by the common denominator, you will not have to deal with fractions.

OpenStudy (texaschic101):

because a division problem is nothing but a fraction

OpenStudy (anonymous):

ok